### v1

@@ 3.2 Batch Reasoning
To conduct the batch reasoning, we guide LLM to predict the stance likelihood $P_{\text{LLM},i}$ and output corresponding explanation $e_{i}$ for each tweet $x_{i}$ in the text batch =$\mathcal{B}=\{(x_{0},t_{0}),(x_{1},t_{1}),...(x_{B},t_{B})\}$:
@@ 3.4 Reasoning-Augmented Training
$\small\ell_{c}(\textbf{h}_{x_{i}})=\log\frac{\sum_{(x_{i},x_{j}^{+})\in\mathcal{P}_{i}}\exp(\text{sim}({\textbf{h}_{x_{i}}},{\textbf{h}_{x_{j}}^{+}})/\tau_{s})}{\sum_{x_{j}\in\mathcal{B}\backslash{x_{i}}}\exp(\text{sim}({\textbf{h}_{x_{i}}},{\textbf{h}_{x_{j}}})/\tau_{s})},$ (12)
where $\lambda$ is weight hyperparameter, $\tau_{s}$ is temperature hyperparameter, $\textbf{h}_{(\cdot)}$ is the embedding of tweet output by SLM. =$\mathcal{B}=\{x_{0},x_{1},...\}$ is mini-batch training data, and \$\mathcal{B}\backslash{x_{i}}$ is $\mathcal{B}$ excluding sample $x_{i}$. $\mathcal{P}_{i}=\{({{x_{i}}},{{x_{0}}^{+}}),({{x_{i}}},{{x_{1}}^{+}}),...\}$ is the set of positive pairs for $x_{i}$, which consists of samples with the same label in mini-batch.
@@ 3.5 Workflow of CoVer
Output: Stance labels =$\mathcal{L}=\{l_{i}\}_{i=1}^{N}$
@@ 4.2.1 Baseline Comparison (RQ1)
Table 1: Zero-Shot Stance Detection Experiments on Sem16 and VAST datasets. The best results are in bold and the second-best results are in underline. Results with * denote that CoVer significantly outperforms baselines with the p-value $<$ 0.05.
@@
[…]
@@
Table 2: Zero-Shot Stance Detection Experiments on the P-stance dataset. The best results are in bold and the second-best results are in underline. Results with * denote that CoVer significantly outperforms baselines with the p-value $<$ 0.05.

### v2

@@ 3.2 Batch Reasoning
To conduct the batch reasoning, we guide LLM to predict the stance likelihood $P_{\text{LLM},i}$ and output corresponding explanation $e_{i}$ for each tweet $x_{i}$ in the text batch $\mathcal{B}=\{(x_{0},t_{0}),(x_{1},t_{1}),...(x_{B},t_{B})\}$:
@@ 3.4 Reasoning-Augmented Training
[…]
@@ 3.5 Workflow of CoVer
Output: Stance labels $\mathcal{L}=\{l_{i}\}_{i=1}^{N}$
@@ 4.2.1 Baseline Comparison (RQ1)
Table 1: Zero-Shot Stance Detection Experiments on Sem16 and VAST datasets. The best results are in bold and the second-best results are in underline. Results with * denote that CoVer significantly outperforms baselines with the p-value << 0.05.
@@
Small-Large Language Model Based Method
@@
Table 2: Zero-Shot Stance Detection Experiments on the P-stance dataset. The best results are in bold and the second-best results are in underline. Results with * denote that CoVer significantly outperforms baselines with the p-value << 0.05.